OpenStudy (anonymous):

A scientist counts 35 bacteria present in a culture and finds that the number of bacteria triples each hour. The function y = 35 • 3x models the number of bacteria after x hours. Graph the function. (You do not need to submit the graph). Use the graph to estimate when there will be about 550 bacteria in the culture.

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

About 2hrs and 20mins

OpenStudy (anonymous):

can u explain in a other way

OpenStudy (anonymous):

@ktnguyen

OpenStudy (anonymous):

the function should be \(A(x)=35\times 3^x\)

OpenStudy (anonymous):

solve the equation \[550=35\times 3^x\] by first dividing by 35 to get \[\frac{110}{7}=3^x\] then change of base to get \[x=\frac{\ln(\frac{110}{7})}{\ln(3)}\]

OpenStudy (anonymous):

i get about 2.5 so two and a half hours if you have not discussed logarithms yet, take out a calculator and try \(35\times 3^x\) with various values for \(x\) when you get to \(x=2.5\) you will see the result is about 550

OpenStudy (anonymous):

thank u thank u